Consider the following information:

Rate of return if state occurs

State of economy

Probability of state of economy

Stock A

Stock B

Boom

0.2

24%

45%

Good

0.35

9%

10%

Poor

0.3

3%

-10%

Bust

??

-5%

-25%

You have $2,000 invested in stock A and $3,000 invested in stock B. Compute the expected return and total risk of this portfolio.
Answer #1
Total Portfolio value = Value of Stock A + Value of Stock B
=2000+3000
=5000
Weight of Stock A = Value of Stock A/Total Portfolio Value
= 2000/5000
=0.4
Weight of Stock B = Value of Stock B/Total Portfolio Value
= 3000/5000
=0.6
Stock A
Scenario Probability Return% =rate of return% * probability Actual return -expected return(A)% (A)^2* probability
Boom 0.2 24 4.8 15.9 0.0050562
Good 0.35 9 3.15 0.9 2.835E-05
Poor 0.3 3 0.9 -5.1 0.0007803
Bust 0.15 -5 -0.75 -13.1 0.00257415
Expected return %= sum of weighted return = 8.1 Sum=Variance Stock A= 0.00844
Standard deviation of Stock A% =(Variance)^(1/2) 9.19
Stock B
Scenario Probability Return% =rate of return% * probability Actual return -expected return(A)% (B)^2* probability
Boom 0.2 45 9 39.25 0.03081125
Good 0.35 10 3.5 4.25 0.000632188
Poor 0.3 -10 -3 -15.75 0.007441875
Bust 0.15 -25 -3.75 -30.75 0.014183438
Expected return %= sum of weighted return = 5.75 Sum=Variance Stock B= 0.05307
Standard deviation of Stock B% =(Variance)^(1/2) 23.04
Covariance Stock A Stock B:
Scenario Probability Actual return% -expected return% for A(A) Actual return% -expected return% For B(B) (A)*(B)*probability
Boom 0.2 15.9 39.25 0.0124815
Good 0.35 0.9 4.25 0.000133875
Poor 0.3 -5.1 -15.75 0.00240975
Bust 0.15 -13.1 -30.75 0.006042375
Covariance=sum= 0.0210675
Correlation A&B= Covariance/(std devA*std devB)= 0.995515623
Expected return%= Wt Stock A*Return Stock A+Wt Stock B*Return Stock B
Expected return%= 0.4*8.1+0.6*5.75
Expected return%= 6.69
Variance =( w2A*σ2(RA) + w2B*σ2(RB) + 2*(wA)*(wB)*Cor(RA, RB)*σ(RA)*σ(RB))
Variance =0.4^2*0.09186^2+0.6^2*0.23037^2+2*0.4*0.6*0.09186*0.23037*0.99552
Variance 0.03057
Standard deviation= (variance)^0.5
Standard deviation= 17.48% = risk